Maximize 4x + 6y subject to the following constraints (using the simplex method)

x + 4y <= 4

3x + 2y <= 6

x>= 0, y>= 0
Answer #1

Print This Solution Close This Solution

Find solution using Simplex(BigM) method
MAX Z = 4x1 + 6x2
subject to
x1 + 4x2 <= 4
3x1 + 2x2 <= 6
and x1,x2 >= 0


Solution:
Problem is

Max Z = 4 x1 + 6 x2
subject to
x1 + 4 x2 4
3 x1 + 2 x2 6
and x1,x2≥0;



The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate

1. As the constraint 1 is of type '≤' we should add slack variable S1

2. As the constraint 2 is of type '≤' we should add slack variable S2

After introducing slack variables

Max Z = 4 x1 + 6 x2 + 0 S1 + 0 S2
subject to
x1 + 4 x2 + S1 = 4
3 x1 + 2 x2 + S2 = 6
and x1,x2,S1,S2≥0


Iteration-1 Cj 4 6 0 0
B CB XB x1 x2 Entering variable S1 S2 MinRatio
XBx2
S1 Leaving variable 0 4 1 (4) (pivot element) 1 0 44=1
S2 0 6 3 2 0 1 62=3
Z=0 0=
Zj=∑CBXB
Zj Zj=∑CBxj 0 0=0×1+0×3
Zj=∑CBx1
0 0=0×4+0×2
Zj=∑CBx2
0 0=0×1+0×0
Zj=∑CBS1
0 0=0×0+0×1
Zj=∑CBS2
Cj-Zj 4 4=4-0 6 6=6-0↑ 0 0=0-0 0 0=0-0



Positive maximum Cj-Zj is 6 and its column index is 2. So, the entering variable is x2.

Minimum ratio is 1 and its row index is 1. So, the leaving basis variable is S1.

∴ The pivot element is 4.

Entering =x2, Departing =S1, Key Element =4

R1(new)=R1(old)÷4

R2(new)=R2(old)-2R1(new)

Iteration-2 Cj 4 6 0 0
B CB XB x1 Entering variable x2 S1 S2 MinRatio
XBx1
x2 6 1 1=4÷4
R1(new)=R1(old)÷4
0.25 0.25=1÷4
R1(new)=R1(old)÷4
1 1=4÷4
R1(new)=R1(old)÷4
0.25 0.25=1÷4
R1(new)=R1(old)÷4
0 0=0÷4
R1(new)=R1(old)÷4
10.25=4
S2 Leaving variable 0 4 4=6-2×1
R2(new)=R2(old)-2R1(new)
(2.5) 2.5=3-2×0.25 (pivot element)
R2(new)=R2(old)-2R1(new)
0 0=2-2×1
R2(new)=R2(old)-2R1(new)
-0.5 -0.5=0-2×0.25
R2(new)=R2(old)-2R1(new)
1 1=1-2×0
R2(new)=R2(old)-2R1(new)
42.5=1.6
Z=6 6=6×1
Zj=∑CBXB
Zj Zj=∑CBxj 1.5 1.5=6×0.25+0×2.5
Zj=∑CBx1
6 6=6×1+0×0
Zj=∑CBx2
1.5 1.5=6×0.25+0×(-0.5)
Zj=∑CBS1
0 0=6×0+0×1
Zj=∑CBS2
Cj-Zj 2.5 2.5=4-1.5↑ 0 0=6-6 -1.5 -1.5=0-1.5 0 0=0-0



Positive maximum Cj-Zj is 2.5 and its column index is 1. So, the entering variable is x1.

Minimum ratio is 1.6 and its row index is 2. So, the leaving basis variable is S2.

∴ The pivot element is 2.5.

Entering =x1, Departing =S2, Key Element =2.5

R2(new)=R2(old)×0.4

R1(new)=R1(old)-0.25R2(new)

Iteration-3 Cj 4 6 0 0
B CB XB x1 x2 S1 S2 MinRatio
x2 6 0.6 0.6=1-0.25×1.6
R1(new)=R1(old)-0.25R2(new)
0 0=0.25-0.25×1
R1(new)=R1(old)-0.25R2(new)
1 1=1-0.25×0
R1(new)=R1(old)-0.25R2(new)
0.3 0.3=0.25-0.25×(-0.2)
R1(new)=R1(old)-0.25R2(new)
-0.1 -0.1=0-0.25×0.4
R1(new)=R1(old)-0.25R2(new)
x1 4 1.6 1.6=4×0.4
R2(new)=R2(old)×0.4
1 1=2.5×0.4
R2(new)=R2(old)×0.4
0 0=0×0.4
R2(new)=R2(old)×0.4
-0.2 -0.2=(-0.5)×0.4
R2(new)=R2(old)×0.4
0.4 0.4=1×0.4
R2(new)=R2(old)×0.4
Z=10 10=6×0.6+4×1.6
Zj=∑CBXB
Zj Zj=∑CBxj 4 4=6×0+4×1
Zj=∑CBx1
6 6=6×1+4×0
Zj=∑CBx2
1 1=6×0.3+4×(-0.2)
Zj=∑CBS1
1 1=6×(-0.1)+4×0.4
Zj=∑CBS2
Cj-Zj 0 0=4-4 0 0=6-6 -1 -1=0-1 -1 -1=0-1



Since all Cj-Zj≤0

Hence, optimal solution is arrived with value of variables as :
x1=1.6,x2=0.6

Max Z=10